OpenStudy (anonymous):

pls help (x^-3)^-5x^6

OpenStudy (anonymous):

answer choices x^21 x^-21 x^48 x^-48

OpenStudy (nurali):

\[(x ^{-3})^{-5}*x^6\] \[\frac{ x^6 }{( \frac{ 1 }{ x^3 })^5}\] \[\frac{ x^6 }{ \frac{ 1 }{ x ^{15} } }\] \[x^6*x ^{15}\] \[x ^{21}\]

OpenStudy (anonymous):

thank you so much

OpenStudy (nurali):

My Pleasure.